Question 12 Is the following trigonometric identity True or False sec(x) sec() sin2 (x) = cos(x) = COS True False

Is the following trigonometric identity True or False
Answer #1

sint (x)+105(x)=1 / Sec (x) sec(a). sin(x) • secas (1 - sit(a)] = sec (2) [ cos(ay] - celom. cos*(a) = 603 () So, [TRUE cos(